# Constraints from CMB lensing tomography with projected bispectra

### v1

@@ 2.5 Inference pipeline
$-2\log p(\mathbf{d}|\vec{\theta})=(\mathbf{d}-\mathbf{t}(\vec{\theta}))^{\mathsf{T}}\mathsf{C}^{-1}(\mathbf{d}-\mathbf{t}(\vec{\theta}))+{\cal K}\,.$ (44)
Here, $\mathbf{d}$ is the data vector containing our combined measurements of the power spectra and bispectra, and $\mathsf{C}$ is the covariance matrix of those measurements, which is obtained as described in sections 2.1.3, 2.2, and Appendix C. The theory predictions $\mathbf{t}(\vec{\theta})$ are generated for a set of parameters $\vec{\theta}$ as described in Section 2.4, and ${\cal K}$ is an irrelevant normalisation constant, independent of $\vec{\theta}$.
The parameter set in our study is relatively small, containing only $\sigma_{8}$, $\Omega_{m}$, and the linear bias $b_{1}$. All other cosmological parameters are simply fixed to the best-fit values found by Planck Collaboration et al. (2020a). The higher-order bias parameters, $b_{2}$ and $b_{s}$ will be determined in terms of $b_{1}$: in our fiducial analysis, we will make use of the empirical relations between these parameters and $b_{1}$ found in Lazeyras et al. (2016) and Lazeyras & Schmidt (2018). We will refer to these as the “coevolution relations” in what follows. These relations have been found to accurately describe the clustering of dark matter haloes as well as galaxies – although with a much larger scatter in the latter case (Barreira et al., 2021). Our purpose with these simplifications is to test the constraining power of the projected galaxy bispectrum through a simple cosmological analysis. We leave a more detailed analysis, in which the higher-order bias parameters are determined self-consistently by including their contribution to the one-loop power spectra, and which incorporates external data to place priors on the cosmological parameters we fix here, for future work.
@@
    To account for the dependence on $\Omega_{m}$ we simply repeat the previous two steps for $100$ different values of this parameter, spanning the range $\Omega_{m}\in[0.05,0.5]$, and generate a cubic spline interpolator for each FSB component as a function of $\Omega_{m}$.
@@
    To generate predictions for a cosmology with a different value of $\sigma_{8}$ and $\Omega_{m}$, we simply rescale each of the templates described above, evaluated at the corresponding value of $\Omega_{m}$, by appropriate powers of the ratio $r_{\sigma_{8}}\equiv\sigma_{8}/\sigma_{8}^{\rm fid}$, where $\sigma_{8}$ is the sampled value of this parameter, and $\sigma_{8}^{\mathrm{fid}}$ is the fiducial value (given in Table 1). The $ggg$ and $gg\kappa$ bispectra both scale with $r_{\sigma_{8}}^{4}$, with the exception of the shot noise terms proportional to the power spectrum, which are scaled as $r_{\sigma_{8}}^{2}$, and the pure noise term in Eq. 33 which is independent of $\sigma_{8}$. In turn, all terms entering the $gg$ and $g\kappa$ power spectra scale with $r_{\sigma_{8}}^{2}$.
@@
[…]
@@
As described in Eq. 44, we approximate the likelihood of the data vector $\mathbf{d}$ to a multivariate normal distribution, described by the data vector covariance $\mathsf{C}$. This is usually a reasonable approximation at high $\ell$ where the Central Limit Theorem (CLT) applies, but can fail at the largest scales; we therefore apply another fiducial $-\ell$ cut at $\ell_{\mathrm{min}}=20$ for all measured observables – $\Phi^{ggg}_{LL\ell}$, $\Phi^{gg\kappa}_{LL\ell}$, $C_{\ell}^{gg}$ and $C_{\ell}^{g\kappa}$. H25 showed that $\Phi^{ggg}_{LL\ell}$ and $C_{\ell}^{gg}$ in fact follow a near-Gaussian distribution, even when including lower multipoles ($\ell>2$).
@@ 2.6 Estimating detection significance
where $\mathbf{x}$ is one FSB measurement, $\bm{\mu}$ is the expectation value of the estimator for a null detection – in our case, the noise terms given by either Eq. 33 or 34. The covariance matrix $\mathsf{C}$ is calculated as described in Section 2.1.3. $N_{\mathrm{dof}}$ is the number of degrees of freedom, in this case the number of data points in $\mathbf{x}$, and represents the expectation value of the previous term in the case of pure noise fluctuations. This method constitutes a good first estimate of the SNR, in that it does not depend on any specific model for the signal we are after. When the SNR is sufficiently high, the argument of the square root is always positive, as $\chi^{2}_{0}\gg N_{\rm dof}$. In low-SNR scenarios, noise fluctuations may make the argument of the square root negative. In such cases, we will simply report ${\rm SNR}=0$. We will refer to this first model-independent estimate of the SNR as “${\rm SNR}_{1}$” in what follows.
@@
Note that here ${\bf T}$ contains only the bispectrum signal (i.e. the FSB version of Eq. 2.4.3 or 2.4.3), with the noise components absorbed in $\bm{\mu}$, as in Eq. 45. We will refer to this second SNR estimate as “${\rm SNR}_{2}$” in what follows. It is worth noting that this SNR estimate is only as accurate as the model used to construct ${\bf T}$ which, as we discussed in the previous section, may not be the case on small scales ($\ell\geq\ell_{\rm max}^{B}$).
@@ 3 Data
3.1 DESI LRGs
The sample is divided into four different redshift bins, covering the redshift range $0.4\lesssim z\lesssim 1$, with number densities in the range $\bar{n}\sim 80$-$160\,{\rm deg}^{-2}$. We use the maps of the LRG overdensity in each redshift bin used in White et al. (2022), and made publicly available [^6] https://zenodo.org/records/5834378 together with the analysis mask. We note that changes were made to the sample used in White et al. (2022), as described in Zhou et al. (2023b), and this revised sample was then used in the cosmological analysis of Sailer et al. (2025b). As explained in Zhou et al. (2023b), the main changes made involve an improvement in the redshift distributions, modifications to the zero-point offsets in the southern sky, and changes in the sky contaminant templates used. This results in a small change in the LRG auto-correlations ($<2\%$). We have verified that none of our FSB measurements (neither $\Phi^{ggg}_{LL\ell}$ nor $\Phi^{gg\kappa}_{LL\ell}$) change significantly within the measurement uncertainties when using the newer sample. Hence, none of the results presented here, regarding the detectability and constraining power of the projected bispectrum, should be affected by these changes.
@@ 4.1 Measurement and detection
Figure 2: The $ggg$ FSBs measured in all four redshift bins. Each column corresponds to one redshift bin, and each row to one of the four colourful filters shown in Fig. 1 – here represented by the greyed-out band. The FSB measurements are shown in blue, and the black solid line shows the theoretical fit for published values of $b_{1}$ (White et al., 2022) and for a fiducial cosmology (Table 1). These theory predictions are used as templates in Eq. 46 to report the $\text{SNR}_{2}$ shown on the colour bars on the right of each FSB. Beyond the conservative cuts used in our inference pipeline (see Section 4.2), we show the predictions as dashed lines to emphasise that the tree-level bispectrum is not necessarily a good theoretical description of the power at those scales. It is worth noting that the FSBs in all but the first filter (i.e. the first row in this plot) also lie in this small-scale regime. The blue shaded area represents the amount of noise (Eq. 33) contributing to the overall theory prediction. Note that we plot $\ell\Phi^{ggg}_{LL\ell}$ instead of $\Phi^{ggg}_{LL\ell}$ to improve readability at higher multipoles.
@@
The naive theoretical expectation, with no free parameters, provides a reasonable description of the $ggg$ FSBs for all except in the first redshift bin, where it systematically underpredicts the signal. As we discuss in Section 4.2.2, this can be accommodated by the combination of a higher value of the linear bias and a lower $\sigma_{8}$ (the latter in qualitative agreement with the lower clustering amplitude seemingly preferred by this sample in White et al., 2022). Interestingly, the theoretical predictions are surprisingly accurate even beyond the scale cuts we will implement in the next section, over which the tree-level description of the galaxy bispectrum has been found to be reliable in 3D analyses (Ivanov et al., 2022). The theoretical predictions are shown as dashed lines on the angular scales outside of the fiducial scale cuts used in Section 4.2. It would be interesting to determine the range of scales over which the projected bispectrum can be safely predicted. Since the projected galaxy overdensity is necessarily more Gaussian than its 3D counterpart, it may be possible to extend the analysis to smaller angular scales, thereby recovering more cosmological information. We leave this study for future work. For now we will remain cautious and implement conservative scale cuts for the analysis in Section 4.2. However, to quantify the significance with which the bispectrum signal is detected in these measurements, independently of our ability to accurately model it, we will use the full range of scales $\ell<2N_{\rm side}$. Visually, the detection of the non-Gaussian contributions to the FSB can be seen by comparing our measurements with the blue band showing the expected contribution from shot noise.
The SNR with which the $ggg$ and $gg\kappa$ FSBs are detected in each redshift bin, combining the signal from all FSB filters, are reported in Table 3. The combined SNR is computed following the two methods described in Section 2.6, using as input the concatenated data vector containing the FSB measurements in all filters. We use the theoretical predictions shown in Figs. 2 and 3 (or rather their pure signal contributions) to build the theoretical template used to estimate ${\rm SNR}_{2}$. It is worth noting that, since the theoretical template is not accurate on small scales (this is clearly visible in the figures as the FSB filter moves to higher $\ell$s and higher filters – as seen also in H25), the approach used to estimate ${\rm SNR}_{2}$ may become unreliable. Thus, we also provide the estimates of ${\rm SNR}_{1}$, which do not rely on any theoretical template (but instead become unreliable for low-SNR data).
@@
The detection of the $ggg$ FSB is indisputable, with a combined $\mathrm{SNR}>30$ in all redshift bins. We find that both estimates, ${\rm SNR}_{1}$ and ${\rm SNR}_{2}$ are in qualitative agreement, with ${\rm SNR}_{1}$ providing a consistently larger estimate (by $\sim 10$-$20\%$). This is likely due to the fact that the model used to create the theoretical template ${\bf T}$ used to calculate ${\rm SNR}_{2}$ (see Eq. 46) under-predicts the signal on small scales, leading to a lower inferred signal amplitude.
In turn, the SNR of the $gg\kappa$ FSBs is significantly lower. Nevertheless, the signal is clearly detected, with ${\rm SNR}>5$ in the three last redshift bins, and only tentatively in the first bin (${\rm SNR}\sim 3$). Judging from Fig. 3, much of the signal seems to be concentrated in the $-\ell$ filters, which will not be used in the cosmological analysis presented in the next section. Thus, although detected, we do not expect $\Phi^{gg\kappa}_{LL\ell}$ to contribute significantly to the final cosmological constraints.
@@ 4.2.1 FSB measurements
As described in Section 2.5, we implement conservative scale cuts to the data to ensure our theoretical predictions are reliable. This means we must also make sure the FSB filters chosen are such that all triangle configurations contributing to the estimator satisfy these cuts. As described in Section 2.3, to ensure this, we use new measurements of the FSB, using a single filter lying below the most conservative $\ell_{\mathrm{max}}^{B}$ cut (corresponding to the lowest redshift bin $z_{1}$, shown in Fig. 1), and reduce the bandpower width to $\Delta\ell=9$. Our data vector, for a given redshift bin, will therefore consist of a single $\Phi^{ggg}_{LL\ell}$ and $\Phi^{gg\kappa}_{LL\ell}$ for this filter, in addition to the power spectrum measurements $C_{\ell}^{gg}$ and $C_{\ell}^{g\kappa}$. These measurements are shown in Fig. 4 for the four redshift bins, truncated at their respective $\ell_{\mathrm{max}}$ cuts. Unsurprisingly, as discussed in the previous section, the SNR of the $\Phi^{gg\kappa}_{LL\ell}$ measurements within these scales is very low (below 1 in all cases). We will therefore discard the $gg\kappa$ measurements in our analysis, keeping only the $ggg$ FSBs, which are detected above $6\sigma$ in all cases. Increasing the SNR should allow us to make the most of the constraining power of the $gg\kappa$ FSBs. This could be achieved by either exploiting all scales within the studied multipole range (allowing for the filter to be defined up to $\ell_{\mathrm{max}}^{B}(z)$ in each bin, instead of using the same, large-scale filter for all) or pushing our model to smaller scales (thus allowing us to relax the conservative scale cuts used so far in this study). Using more sensitive CMB lensing data, from ground-based experiments such as the Atacama Cosmology Telescope (Madhavacheril et al., 2024), could also improve the SNR of these measurements. We will study these potential avenues for optimisation in future work.

### v2

@@ 2.5 Inference pipeline
$2\log p(\mathbf{d}|\vec{\theta})=(\mathbf{d}-\mathbf{t}(\vec{\theta}))^{\mathsf{T}}\mathsf{C}^{-1}(\mathbf{d}-\mathbf{t}(\vec{\theta}))+{\cal K}\,.$ (44)
Here, $\mathbf{d}$ is the data vector containing our combined measurements of the power spectra and bispectra (however we will see in Section 4.2 that the $gg\kappa$ bispectrum will ultimately not be included in the data vector), and $\mathsf{C}$ is the covariance matrix of those measurements, which is obtained as described in sections 2.1.3, 2.2, and Appendix C. The theory predictions $\mathbf{t}(\vec{\theta})$ are generated for a set of parameters $\vec{\theta}$ as described in Section 2.4, and $\cal K$} is an irrelevant normalisation constant, independent of $\vec{\theta}$.
The parameter set in our study is relatively small, containing only $\sigma_{8}$, $\Omega_{m}$, and the linear bias $b_{1}$. All other cosmological parameters are simply fixed to the best-fit values found by Planck Collaboration et al. (2020a). The higher-order bias parameters, $b_{2}$ and $b_{s}$ will be determined in terms of $b_{1}$: in our fiducial analysis, we will make use of the empirical relations between these parameters and $b_{1}$ found in Lazeyras et al. (2016) and Lazeyras and Schmidt (2018). We will refer to these as the “coevolution relations” in what follows. These relations have been found to accurately describe the clustering of dark matter haloes as well as galaxies – although with a much larger scatter in the latter case (Barreira et al., 2021). Our purpose with these simplifications is to test the constraining power of the projected galaxy bispectrum through a simple cosmological analysis. We leave a more detailed analysis, in which the higher-order bias parameters are determined self-consistently by including their contribution to the one-loop power spectra, and which incorporates external data to place priors on the cosmological parameters we fix here, for future work.
@@
    To account for the dependence on $\Omega_{m}$ we simply repeat the previous two steps for $100$ different values of this parameter, spanning the range $\Omega_{m}\in[0.05,0.5]$, and generate a cubic spline interpolator for each FSB component as a function of $\Omega_{m}$. We provide an overview of the final interpolator performance, both in terms of runtime and accuracy, in Appendix D.
@@
    To generate predictions for a cosmology with a different value of $\sigma_{8}$ and $\Omega_{m}$, we simply rescale each of the templates described above, evaluated at the corresponding value of $\Omega_{m}$, by appropriate powers of the ratio $r_{\sigma_{8}}\equiv\sigma_{8}/\sigma_{8}^{\rm fid}$, where $\sigma_{8}$ is the sampled value of this parameter, and $\sigma_{8}^{\mathrm{fid}}$ is the fiducial value (given in Table 1). The gggggg and $gg\kappa$ bispectra both scale with $r_{\sigma_{8}}^{4}$, with the exception of the shot noise terms proportional to the power spectrum, which are scaled as $r_{\sigma_{8}}^{2}$, and the pure noise term in Eq. 33 which is independent of $\sigma_{8}$. In turn, all terms entering the $gg$ and $g\kappa$ power spectra scale with $r_{\sigma_{8}}^{2}$.
@@
To ensure that our theoretical model is valid, we introduce scales cuts in the data vector: we only select scales below $\ell_{\mathrm{max}}=\chi(\bar{z})k_{\mathrm{max}}$, where $\bar{z}$ is the effective redshift of the galaxy sample. The value of $k_{\mathrm{max}}$ depends on the observable: we set $k_{\mathrm{max}}^{P}=0.07\,\mathrm{Mpc}^{-1}$ for the power spectrum, and a more conservative $k_{\mathrm{max}}^{B}=0.05\,\mathrm{Mpc}^{-1}$ for the bispectrum. The value of $k_{\rm max}^{P}$ matches that used by Sailer et al. (2025b) in the analysis of the LRG sample used here assuming a linear bias relation. Although the tree-level matter bispectrum usually provides accurate fits to simulation results for scales up to $k_{\mathrm{max}}\lesssim 0.1\,h{\rm Mpc}^{-1}$ (see e.g. Takahashi et al., 2020), our $k_{\mathrm{max}}^{B}$ cuts are motivated by Ivanov et al. (2022), where a scale cut $k_{\mathrm{max}}=0.08\,h\mathrm{Mpc}^{-1}$ is used to model the tree-level bispectrum model in redshift space at $z=0.61$ (which lies between the first and second redshift bins used here, see Section 3.1). For our fiducial cosmology described in Table 1, this cut corresponds to about $k_{\mathrm{max}}^{B}=0.05\,\mathrm{Mpc}^{-1}$, which in turn gives $\ell_{\mathrm{max}}^{P}=130,\,166,\,200,\,226$ and $\ell_{\mathrm{max}}^{B}=93,\,119,\,143,\,161$ in the four redshift bins studied here. We do not expect these conservative cuts to vary greatly for similar values of $\sigma_{8}$ or $h$. Note that this conservative $k_{\mathrm{max}}^{B}$ cut was necessary in Ivanov et al. (2022) due to the presence of Redshift-Space Distortions (RSDs), which do not affect 2D projected probes as much as their 3D counterparts. It is likely that this scale cut is thus conservative for our purposes, and determining the smallest physical scale up to which projected bispectra may be described by simple perturbation theory models would allow us to improve upon any constraints derived in this analysis.
@@
As described in Eq. 44, we approximate the likelihood of the data vector $\mathbf{d}$ to a multivariate normal distribution, described by the data vector covariance $\mathsf{C}$. This is usually a reasonable approximation at high $\ell$ where the Central Limit Theorem (CLT) applies, but can fail at the largest scales; we therefore apply another fiducial low-$\ell$ cut at $\ell_{\mathrm{min}}=20$ for all measured observables – $\Phi^{ggg}_{LL\ell}$, $\Phi^{gg\kappa}_{LL\ell}$, $C_{\ell}^{gg}$ and $C_{\ell}^{g\kappa}$. This should also help mitigate any large-scale inaccuracies introduced by the Limber approximation and neglecting RSD contributions to both FSBs – although the filter used to define the FSB does include large scales, below $\ell_{\rm min}$, in the bispectrum. As shown in Verdiani et al. (2025), however, the larger statistical uncertainties in this statistic makes these effects negligible. H25 furthermore showed that $\Phi^{ggg}_{LL\ell}$ and $C_{\ell}^{gg}$ in fact follow a near-Gaussian distribution, even for bandpowers including lower multipoles ($\ell>2$).
@@ 2.6 Estimating detection significance
where $\mathbf{x}$ is one FSB measurement, $\bm{\mu}$ is the expectation value of the estimator for a null detection – in our case, the noise terms given by either Eq. 33 or 34. The covariance matrix $\mathsf{C}$ is calculated as described in Section 2.1.3. $N_{\mathrm{dof}}$ is the number of degrees of freedom, in this case the number of data points in $\mathbf{x}$, and represents the expectation value of the previous term in the case of pure noise fluctuations. This method constitutes a good first estimate of the SNR, in that it does not depend on any specific model for the signal we are after. When the SNR is sufficiently high, the argument of the square root is always positive, as $\chi^{2}_{0}\gg N_{\rm dof}$. In low-SNR scenarios, noise fluctuations may make the argument of the square root negative. In such cases, we will simply report ${\rm SNR}=0$. We will refer to this first model-independent estimate of the SNR as ${\rm SNR}_{1}$” in what follows.
@@
Note that here $\bf T$} contains only the bispectrum signal (i.e. the FSB version of Eq. 2.4.3 or 2.4.3), with the noise components absorbed in $\bm{\mu}$, as in Eq. 45. We will refer to this second SNR estimate as ${\rm SNR}_{2}$” in what follows. It is worth noting that this SNR estimate is only as accurate as the model used to construct $\bf T$} which, as we discussed in the previous section, may not be the case on small scales ($\ell\geq\ell_{\rm max}^{B}$).
@@ 3 Data
3.1 DESI LRGs
The sample is divided into four different redshift bins, covering the redshift range $0.4\lesssim z\lesssim 1$, with number densities in the range $\bar{n}\sim 80$-$160\,{\rm deg}^{-2}$. We use the maps of the LRG overdensity in each redshift bin used in White et al. (2022), and made publicly available [^7] https://zenodo.org/records/5834378 together with the analysis mask. We note that changes were made to the sample used in White et al. (2022), as described in Zhou et al. (2023b), and this revised sample was then used in the cosmological analysis of Sailer et al. (2025b). As explained in Zhou et al. (2023b), the main changes made involve an improvement in the redshift distributions, modifications to the zero-point offsets in the southern sky, and changes in the sky contaminant templates used. This results in a small change in the LRG auto-correlations ($2\%$). We have verified that none of our FSB measurements (neither $\Phi^{ggg}_{LL\ell}$ nor $\Phi^{gg\kappa}_{LL\ell}$) change significantly within the measurement uncertainties when using the newer sample. Hence, none of the results presented here, regarding the detectability and constraining power of the projected bispectrum, should be affected by these changes.
@@ 4.1 Measurement and detection
Figure 2: The gggggg FSBs measured in all four redshift bins. Each column corresponds to one redshift bin, and each row to one of the four colourful filters shown in Fig. 1 – here represented by the greyed-out band. The FSB measurements are shown in blue, and the black solid line shows the theoretical fit for published values of $b_{1}$ (White et al., 2022) and for a fiducial cosmology (Table 1). These theory predictions are used as templates in Eq. 46 to report the $\text{SNR}_{2}$ shown on the colour bars on the right of each FSB. Beyond the conservative cuts used in our inference pipeline (see Section 4.2), we show the predictions as dashed lines to emphasise that the tree-level bispectrum is not necessarily a good theoretical description of the power at those scales. It is worth noting that the FSBs in all but the first filter (i.e. the first row in this plot) also lie in this small-scale regime. The blue shaded area represents the amount of noise (Eq. 33) contributing to the overall theory prediction. Note that we plot $\ell\Phi^{ggg}_{LL\ell}$ instead of $\Phi^{ggg}_{LL\ell}$ to improve readability at higher multipoles.
@@
We expect the tree-level description of the galaxy bispectrum to be reliable when all three legs $(\ell_{1},\ell_{2},\ell_{3})$ remain below the fiducial multipole cuts we describe in Section 2.5. This regime corresponds to only a subset of Figs. 2 and 3:
  • Only the first filter selects scales $(\ell_{1},\ell_{2})$ below these scale cuts, whereas the three other filters select smaller scales where linear perturbation does not necessarily hold;
  • We show the predictions as bold dark lines where $\ell_{3}$ (which corresponds to the x-axis) falls below the aforementioned scale cuts, and as dashed lines beyond.
Therefore, our theoretical predictions should be most reliable in the first row of each figure, and where theoretical predictions are shown in bold. The naive theoretical expectation, with no free parameters, provides a reasonable description of the gggggg FSBs within these limits for all except in the first redshift bin, where it systematically underpredicts the signal. As we discuss in Section 4.2.2, this can be accommodated by the combination of a higher value of the linear bias and a lower $\sigma_{8}$ (the latter in qualitative agreement with the lower clustering amplitude seemingly preferred by this sample in White et al., 2022). Interestingly, the theoretical predictions are surprisingly accurate even beyond these scale cuts: in the first and second row of Figs. 2 and 3, the tree-level bispectrum still provides a somewhat reasonable fit to the measurements over an extended multipole range. It would be interesting to determine the range of scales over which the projected bispectrum can be safely predicted. Since the projected galaxy overdensity is necessarily more Gaussian than its 3D counterpart, it may be possible to extend the analysis to smaller angular scales, thereby recovering more cosmological information. We leave this study for future work. For now we will remain cautious and implement conservative scale cuts for the analysis in Section 4.2. However, to quantify the significance with which the bispectrum signal is detected in these measurements, independently of our ability to accurately model it, we will use the full range of scales $\ell<2N_{\rm side}$. Visually, the detection of the non-Gaussian contributions to the FSB can be seen by comparing our measurements with the blue band showing the expected contribution from shot noise.
[…]
@@
The detection of the gggggg FSB is indisputable, with a combined $\mathrm{SNR}>30$ in all redshift bins. We find that both estimates, ${\rm SNR}_{1}$ and ${\rm SNR}_{2}$ are in qualitative agreement, with ${\rm SNR}_{1}$ providing a consistently larger estimate (by $\sim 10$-$20\%$). This is likely due to the fact that the model used to create the theoretical template $\bf T$} used to calculate ${\rm SNR}_{2}$ (see Eq. 46) under-predicts the signal on small scales, leading to a lower inferred signal amplitude.
In turn, the SNR of the $gg\kappa$ FSBs is significantly lower. Nevertheless, the signal is clearly detected, with ${\rm SNR}>5$ in the three last redshift bins, and only tentatively in the first bin (${\rm SNR}\sim 3$). Judging from Fig. 3, much of the signal seems to be concentrated in the higher-$\ell$ filters, which will not be used in the cosmological analysis presented in the next section. Thus, although detected, we do not expect $\Phi^{gg\kappa}_{LL\ell}$ to contribute significantly to the final cosmological constraints.
@@ 4.2.1 FSB measurements
As described in Section 2.5, we implement conservative scale cuts to the data to ensure our theoretical predictions are reliable. This means we must also make sure the FSB filters chosen are such that all triangle configurations contributing to the estimator satisfy these cuts. As described in Section 2.3, to ensure this, we use new measurements of the FSB, using a single filter lying below the most conservative $\ell_{\mathrm{max}}^{B}$ cut (corresponding to the lowest redshift bin $z_{1}$, shown in Fig. 1), and reduce the bandpower width to $\Delta\ell=9$. Our data vector, for a given redshift bin, will therefore consist of a single $\Phi^{ggg}_{LL\ell}$ and $\Phi^{gg\kappa}_{LL\ell}$ for this filter, in addition to the power spectrum measurements $C_{\ell}^{gg}$ and $C_{\ell}^{g\kappa}$. These measurements are shown in Fig. 4 for the four redshift bins, truncated at their respective $\ell_{\mathrm{max}}$ cuts. Unsurprisingly, as discussed in the previous section, the SNR of the $\Phi^{gg\kappa}_{LL\ell}$ measurements within these scales is very low (below 1 in all cases). We will therefore discard the $gg\kappa$ measurements in our analysis, keeping only the gggggg FSBs, which are detected above $6\sigma$ in all cases. Increasing the SNR should allow us to make the most of the constraining power of the $gg\kappa$ FSBs. This could be achieved by either exploiting all scales within the studied multipole range (allowing for the filter to be defined up to $\ell_{\mathrm{max}}^{B}(z)$ in each bin, instead of using the same, large-scale filter for all) or pushing our model to smaller scales (thus allowing us to relax the conservative scale cuts used so far in this study). Using more sensitive CMB lensing data, from ground-based experiments such as the Atacama Cosmology Telescope (Madhavacheril et al., 2024), could also improve the SNR of these measurements. We will study these potential avenues for optimisation in future work.